OpenStudy (anonymous):

whats y= x + 3 8x-7y=12

pooja195 (pooja195):

Are we solving for x and y?

pooja195 (pooja195):

y= x + 3 <---equation 1 8x-7y=12 <-----equation 2 We are given 2 equations we are also given the value of y So we take equation 1 and substitute it into equation 2 like this: \[\huge~\rm~ 8x-7=12+x+3\] from here solve for x :)